A 2 8x10 kg jet plane is ready for takeoff. The jet's engines supply a constant thrust of 6.35x10'N How much runway will it need to reach a minimum takeoff speed of 285km/hr? What is the normal force acting on the jet?


1
Expert's answer
2021-02-01T12:41:30-0500

"285(km\/h)=79.2(m\/s)"


"m=2.8\cdot10^5(kg)"


"F=6.35\cdot 10^6(N)"


"a=F/m=6.35\cdot10^6/(2.8\cdot10^5)=22.7(m/s^2)"


"s=\frac{79.2^2}{2\cdot22.7}=138.1(m)" . Answer


"N=mg=2.8\cdot10^5\cdot 9.8=2.744\cdot 10^6\ (N)" . Answer